OpenStudy (anonymous):

log3(2−x)−1/2log3( x+6)=1

OpenStudy (alfie):

\[1 = \log_33\] Then apply the logs properties.

jhonyy9 (jhonyy9):

(2-x)/sqrt(x+6)=3

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

2-x=3sqrt(x+6)

OpenStudy (anonymous):

can u help me to simplify it by solving for x?

jhonyy9 (jhonyy9):

yes square both sides

jhonyy9 (jhonyy9):

4-4x+x2=9(x+6)

jhonyy9 (jhonyy9):

x2-4x+4=9x+54 x2-13x-50=0

jhonyy9 (jhonyy9):

x1,2=(13+/- sqrt(169+200))/2 =(13+/- sqrt(369))/2